OpenStudy (anonymous):

For the following functions, f and g, determine (a) f(g(x)) (b) g(f(x)) If f(x)= 7x+4 and g(x)-10x+6 , then (a) f(g(x)) (b) g(f(x))

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

it is \(g(x)=10x+6\) ?

OpenStudy (anonymous):

-10x + 6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would you be able to help me with this problem?

OpenStudy (anonymous):

\[f(g(x))=f(-10x+6)=7(-10x+6)+4\] then some algebra to clean up the answer

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

oh so basically we plug them in together equaling to each other?

OpenStudy (anonymous):

oh no not equal to each other

OpenStudy (anonymous):

okay from the basic algebra how do we find f(g(x))?

OpenStudy (anonymous):

look at the steps i wrote first i wrote \(f(g(x))\) then i replaced \(g(x)\) by your \(g(x)\) namely \(-10x+7\)

OpenStudy (anonymous):

so I got this: -70X + 46 = -70X-40

OpenStudy (anonymous):

then evaluate the function \(f\) replacing \(x\) by \(-10x+7\)

OpenStudy (anonymous):

there is no equation to solve here

OpenStudy (anonymous):

-10x+6

OpenStudy (anonymous):

the solution is \[f(g(x))=f(-10x+6)=7(-10x+6)+4=-70x+42+4=-70x+46\]

OpenStudy (anonymous):

right... -70x + ......

OpenStudy (anonymous):

I did the exact same thing. Am i missing something here?

OpenStudy (anonymous):

-70x +46 ....right